OpenStudy (anonymous):

How do I simplify (-27)^-1/3

OpenStudy (anonymous):

same as \[1/\sqrt[3]{-27} which comes \to -3\]

OpenStudy (anonymous):

1/cuberoot(-27) which equals -3

OpenStudy (zzr0ck3r):

1/(-27)^(1/3) = 1/-3 = -1/3